OpenStudy (anonymous):

ASAP!!What is the point-slope equation of a line with slope -5 that contains the point (6, 3)?

OpenStudy (anonymous):

A. y + 3 = -5(x + 6) B. y + 3 = -5(x - 6) C. y - 3 = -5(x - 6) D. y - 3 = -5(x + 6)